Please help me with this problem. I would really appreciate it!
There are 28 students in a class. 20 students have atleast one sibling. 17 have atleast one cousin. Out of those 12 have atleast one sibling and one cousin. how many classmates do not have any siblings or cousins.
the answer choices are: 1, 2, 3, or 4.

Here is how I attempted this:
28 in class.
20 = 1 sibling
17= 1 cousin
12 = 1 cousin and 1 sibling
I thought that there would be 8 students that would be left without anyone, but that wasn't an answer choice. Then I thought that 17-12=5, but also not an answer choice. Please help me. I would really appreciate it! Thank you so much!

Answer by venugopalramana(3286)

SEE MY WORKING BELOW
Please help me with this problem. I would really appreciate it!
There are 28 students in a class.
20 students have atleast one sibling.Out of those 12 have atleast one sibling and one cousin.THAT IS THESE 20 INCLUDE THOSE WITH ATLEAST ONE SIBLING AND THOSE WITH ATLEAST ONE SIBBLING AND ONE COUSIN.HENCE THOSE WITH ONLY ONE SIBLING =20-12=8
17 have atleast one cousin.Out of those 12 have atleast one sibling and one cousin.THAT IS THESE 17 INCLUDE THOSE WITH ATLEAST ONE COUSIN AND THOSE WITH ATLEAST ONE SIBBLING AND ONE COUSIN.HENCE THOSE WITH ONLY ONE COUSIN =17-12=5
how many classmates do not have any siblings or cousins.
SO WE HAVE 8 WITH ONLY SIBLINGS+5 WITH ONLY COUSINS+12 WITH BOTH SIBLING AND COUSINS=TOTAL 25 HAVE SIBLING OR COUSIN OR BOTH
SO THOSE WHO DO NOT HAVE ANY SIBLING OR COUSIN=28-25=3...IS THE ANSWER
the answer choices are: 1, 2, 3, or 4.